OpenStudy (anonymous):

What is the smallest positive value for x such that 180x is a perfect cube?

OpenStudy (dumbcow):

180x = n^3 x = n^3/180 try n=6,7,8... until you get integer for x

OpenStudy (dumbcow):

x=150